OpenStudy (anonymous):

Can someone help me? Porfavor? (: the vertices of square PQRS are P(-4,0) Q(4,3) R(7,-5) S(-1,-8). show that the diagonals of square PQRS are congruent perpendicular bisectors of each other.

OpenStudy (anonymous):

Mid point of P and R is ((-4+7)/2 ,(0-5)/2) = (3/2,-5/2) Mid point of Q and S is ((4-1)/2 ,(3-8)/2) = (3/2,-5/2) So mid point of both is same so they bisect each other. For perpendicular, RP(vector) = (-4-7,0+5) = (-11,5) QS(vector) = (-1-4,-8-3) = (-5,-11) RP.QS = 55-55=0 So Digonals are perpendicular Length of PR =sqrt(11^2 + 5^2) = Length of QS Hence Proved Congruent Perpendicular Bisectors.

OpenStudy (anonymous):

my teacher wants me to prove i t and i've tried to work it..but it just won't come out..